An organic compound was analyzed and found to cont
Answer #1

moles of C = 55.8 / 12 = 4.65

moles of H = 7.03 / 1 = 7.03

moles of O = 37.2 / 16 = 2.325

C                O            H

4.65          2.325     7.03

2                 1                    3

C2H3O -----------------------> empirical formula

empirical formula mass = 24 + 3 + 16 = 43 g/mol

volume = 0.530 L

pressure = 0.974 atm

temperature = 373 K

P V = n R T

0.974 x 0.530 = 1.5 x 0.0821 x 373 / M

M0lar mass = 88.98 g/mol

n = molar mass / empirical formula mass

    = 88.98 / 43

     = 2 (nearly)

molecular formula = n x empirical formula

                             = C4H6O2